OpenStudy (anonymous):

Hey, I need help with one question about ellipses. I have to go so a simple step by step for the situation and then with my numbers to plug in would be good. Please help quickly... The question is: Write the equation of an ellipse with center (-3,2); covertex (-8,2) and focus (-3,-10)

OpenStudy (loser66):

What is the formula of a shifted ellipse?

OpenStudy (anonymous):

\[\frac{ (x-h)^2 }{ a^2 } + \frac{ (y-k)^2 }{ b^2 } = 1\]?

OpenStudy (loser66):

Yup!! how to find the major vertices from foci and co-vertices?

OpenStudy (loser66):

c^2=......

OpenStudy (anonymous):

c^2 = a^2 - b^2

OpenStudy (loser66):

Yes, you have foci c, and b, how to find a^2?

OpenStudy (anonymous):

what what is c? is it -3?

OpenStudy (anonymous):

sorry you kind of lost me...

OpenStudy (loser66):

no no, I am just asking about the formula, not number yet.

OpenStudy (anonymous):

oh ok. I think to find a^2 you turn it around so it is: a^2 = b^2 = c^2? or is it a completely different formula

OpenStudy (loser66):

ok, now, draw out the given points.

OpenStudy (anonymous):



OpenStudy (loser66):